# Hierarchical Gaussian Mixture Normalizing Flow Modeling for Unified Anomaly Detection

### v1

@@ Abstract
Unified anomaly detection (AD) is one of the most challenges for anomaly detection, where one unified model is trained with normal samples from multiple classes with the objective to detect anomalies in these classes. For such a challenging task, popular normalizing flow (NF) based AD methods may fall into a “homogeneous mapping” issue, where the NF-based AD models are biased to generate similar latent representations for both normal and abnormal features, and thereby lead to a high missing rate of anomalies. In this paper, we propose a novel Hierarchical Gaussian mixture normalizing flow modeling method for accomplishing unified Anomaly Detection, which we call HGAD. Our HGAD consists of two key components: inter-class Gaussian mixture modeling and intra-class mixed class centers learning. Compared to the previous NF-based AD methods, the hierarchical Gaussian mixture modeling approach can bring stronger representation capability to the latent space of normalizing flows, so that even complex multi-class distribution can be well represented and learned in the latent space. In this way, we can avoid mapping different class distributions into the same single Gaussian prior, thus effectively avoiding or mitigating the “homogeneous mapping” issue. We further indicate that the more distinguishable different class centers, the more conducive to avoiding the bias issue. Thus, we further propose a mutual information maximization loss for better structuring the latent feature space. We evaluate our method on four real-world AD benchmarks, where we can significantly improve the previous NF-based AD methods and also outperform the SOTA unified AD methods.
@@ 1 Introduction
Anomaly detection has received increasingly wide attentions and applications in different scenarios, such as industrial defect detection [7, 22], video surveillance [1, 30], medical lesion detection [32, 44], and road anomaly detection [36, 10]. Considering the highly scarce anomalies and diverse normal classes, most previous AD studies have mainly devoted to unsupervised one-class learning, i.e., learning one specific AD model by only utilizing one-class normal samples and then detecting anomalies in this class. However, such a one-for-one paradigm would require more human labor, time, and computation costs when training and testing on many product categories, and also underperform when the one normal class has large intra-class diversity.
@@
In this work, we aim to tackle a more practical task: unified anomaly detection. As shown in Fig. 1b, one unified model is trained with normal samples from multiple classes, and the objective is to detect anomalies for all these classes without any fine-tuning. Nonetheless, solving such a task is quite challenging. Currently, there are two reconstruction-based AD methods for tackling the challenging unified AD task, UniAD [40] and PMAD [39]. But the reconstruction-based methods may fall into the “identical shortcut reconstruction” dilemma [40], where anomalies can also be well reconstructed, resulting in the failure of anomaly detection. UniAD and PMAD attempt to mask the adjacent or suspicious anomalies to avoid identical reconstruction. However, due to the diverse scale and shape of anomalies, the masking mechanism cannot completely avoid the abnormal information leakage during reconstruction, the risk of identical reconstruction is still existing. To this end, we consider designing the unified AD model from the normal data distribution learning perspective. The advantage is that we will no longer face the abnormal information leakage risk in principle, as there is no need to reconstruct anomalies as normals for detecting anomalies. Specifically, we employ normalizing flows (NF) to learn the normal data distribution [15].
However, we find that the NF-based AD methods perform unsatisfactorily when applied to the unified AD task. They usually fall into a “homogeneous mapping” issue (see Sec. 3.2), where the NF-based AD models are biased to generate large log-likelihoods for both normal and abnormal inputs (see Fig. 2b). We further explain this issue as: the multi-class distribution is far more diverse and usually multi-modal. However, conventional NF-based AD methods [15, 41] employ the uni-modal Gaussian prior to learn the invertible mapping. This can be seen as learning a mapping from a heterogeneous space to the latent homogeneous space. To learn the mapping well, the network may be prompted to take a bias to concentrate on the coarse-grained common characteristics (e.g., local pixel correlations) and suppress the fine-grained distinguishable characteristics (e.g., semantic content) among different class features [17]. Consequently, the network homogeneously maps different class features to the close latent embeddings. Thus, even anomalies can obtain large log-likelihoods and become less distinguishable.
@@
1. We propose a novel unified AD method: HGAD, where the hierarchical Gaussian mixture modeling approach can bring stronger representation capability to the latent space of normalizing flows for accomplishing the unified anomaly detection task.
@@ 2 Related Work
Anomaly Detection. 1) Reconstruction-based approaches are the most popular AD methods. These methods rely on the assumption that models trained by normal samples would fail in abnormal regions. Many previous works attempt to train AutoEncoders [20, 42], Variational AutoEncoders [18] and GANs [28, 2] to reconstruct the input images. However, these methods face the “identical shortcut” problem [40]. 2) Embedding-based approaches recently show better AD performance by using ImageNet pre-trained networks as feature extractors [6, 11]. PaDiM [12] extract pre-trained features to model Multivariate Gaussian distribution for normal samples, then utilize Mahalanobis distance to measure the anomaly scores. PatchCore [22] extends on this line by utilizing locally aggregated features and introducing greedy coreset subsampling to form nominal feature banks. 3) Knowledge distillation assumes that the student trained to learn the teacher on normal samples could only regress normal features but fail in abnormal features [8]. Recent works mainly focus on feature pyramid [26, 37], reverse distillation [13], and asymmetric distillation [25]. 4) Unified AD approaches attempt to train a unified AD model to accomplish anomaly detection for multiple classes. UniAD [40], PMAD [39] and OmniAL [46] are three existing methods in this new direction. UniAD is a transformer-based reconstruction model with three improvements, it can perform well under the unified case by addressing the “identical shortcut” issue. PMAD is a MAE-based patch-level reconstruction model, which can learn a contextual inference relationship within one image rather than the class-dependent reconstruction mode. OmniAL is a unified CNN framework with anomaly synthesis, reconstruction and localization improvements. To prevent the identical reconstruction, OmniAL trains the model with proposed panel-guided synthetic anomaly data rather than directly using normal data.
@@
Normalizing Flows in Anomaly Detection. In anomaly detection, normalizing flows are employed to learn the normal data distribution [23, 15, 41, 38], which maximize the log-likelihoods of normal samples during training. Rudolph et al. [23] first employ NFs for anomaly detection by estimating the distribution of pre-trained features. In CFLOW-AD [15], the authors further construct NFs on multi-scale feature maps to achieve anomaly localization. Recently, fully convolutional normalizing flows [24, 41] have been proposed to improve the accuracy and efficiency of anomaly detection. In BGAD [38], the authors propose a NF-based AD model to tackle the supervised AD task. In this paper, we mainly propose a novel NF-based AD model (HGAD) with three improvements to achieve much better unified AD performance.
@@ 3.2 Revisiting Normalizing Flow Based Anomaly Detection Methods
Figure 2: Comparison between one-for-one and unified NF-based AD methods on MVTecAD. (a) and (c) show the training losses and the testing anomaly detection and localization AUROCs. (b) shows that the one-for-one NF-based AD model may have an obvious normal-abnormal overlap under the unified case, while ours (d) can bring better normal-abnormal distinguishability.
@@ 3.3 Hierarchical Gaussian Mixture Normalizing Flow Modeling
Figure 3: Model overview. The extracted feature vectors are sent into the normalizing flow model for transforming into latent embeddings. Following [15], we add positional embeddings to each invertible layer as they are effective for NF-based AD methods. Then, we employ our hierarchical Gaussian mixture modeling approach to fit the latent embeddings, which can assist the model against learning the “homogeneous mapping”.
@@ 3.4 Anomaly Scoring
As shown in Alg. 1, we actually extract K-level feature maps and construct the NF model at each level. For each test input feature $x^{k}$ from level-$k$, $k\in K$, we can calculate its intra-class log-likelihood ${\rm log}p_{\theta}(x^{k})=\mathop{{\rm logsumexp}}_{i}(-||\varphi_{\theta}(x^{k})-\mu_{i}^{y}||^{2}_{2}/2+c_{i}^{y})+{\rm log}|{\rm det}J|-d/2{\rm log}(2\pi)$ and inter-class negative entropy $nh(x^{k})=\sum_{y}\mathop{{\rm softmax}}_{y}(-||\varphi_{\theta}(x^{k})-\mu_{1}^{y^{\prime}}||^{2}_{2}/2)\cdot\mathop{{\rm logsoftmax}}_{y}(-||\varphi_{\theta}(x^{k})-\mu_{1}^{y^{\prime}}||^{2}_{2}/2)$. Note that $y$ in $c_{i}^{y}$ denotes which class $x^{k}$ belongs to, not whether $x^{k}$ is normal or abnormal. Next, we convert the log-likelihood to likelihood $p_{\theta}(x^{k})={\rm e}^{{\rm log}p_{\theta}(x^{k})}$. Then, we upsample all $p_{\theta}(x^{k})$ in the level-$k$ to the input image resolution $(H\times W)$ using bilinear interpolation $P_{k}=b(p_{\theta}(x^{k}))\in\mathbb{R}^{H\times W}$. Finally, we calculate anomaly score map $S_{l}$ by aggregating all upsampled likelihoods as $S_{l}=1-\sum_{k=1}^{K}P_{k}$ (see Sec. 3.1). For the inter-class negative entropy, we also follow the above steps to convert to anomaly score map $S_{e}$. Then the final anomaly score map is obtained by combining the two maps $S=S_{l}\odot S_{e}$, where the $\odot$ is the element-wise multiplication. In this way, even if anomalies fall into the inter-class Gaussian mixture distribution, they are usually in the low-density regions among the inter-class class centers. So, we can still ensure that anomalies are out-of-distribution through intra-class log-likelihoods (please see App. 0.A.3 for more discussions).
@@ 3.5 Further Discussions
More Discussions with “identical shortcut”. The “identical shortcut” is essentially caused by the leakage of abnormal information. The process of reconstruction is to remove abnormal information in the input, resulting in the failure of reconstruction in abnormal regions. But if the abnormal features are leaked into the output, this will result in the reconstruction network directly returning a copy of the input as output. This issue usually can be addressed by masking, such as the neighbor masking mechanism in UniAD [40]. However, the “homogeneous mapping” is a specific issue in normalizing flow (NF) based AD methods. In previous NF-based AD methods, the latent feature space is uni-modal. When used for unified anomaly detection, we need to map different class features to the single latent center, this may cause the model more prone to take a bias to map different input features to similar latent features. Thus, with the bias, the log-likelihoods of abnormal features will become closer to the log-likelihoods of normal features, causing normal misdetection or abnormal missing detection. We call this phenomenon as the “homogeneous mapping” issue, rather than casually introducing it. Moreover, as analyzed in Sec. 3.2, we provide a reasonable explanation from the perspective of the formula in normalizing flow. To address this issue, we propose the hierarchical Gaussian Mixture modeling approach, the key designs in our method are completely different from those in UniAD. As the causes and solutions of the two issues are significantly different, “homogeneous mapping” is not intrinsically equal to “identical shortcut”.
In App. 0.A, we also further discuss the limitations, applications, effective guarantee, and complexity of our method. We also provide an information-theoretic view in App. 0.F.
@@ 4.2 Main Results
Setup. For each dataset, we train one unified model with images from different classes. All the images are resized and cropped to $256\times 256$ resolution. The feature maps from stage-1 to stage-3 of EfficientNet-b6 [31] are used as inputs to the normalizing flow models. The parameters of the feature extractor are frozen during training. The layer numbers of the NF models are all 12. The number of inter-class centers is always equal to the number of classes in the dataset. The number of intra-class centers is set as $10$ for all datasets (see ablation study in Sec. 4.3). We use the Adam [21] optimizer with weight decay $1{\rm e}^{-4}$ to train the model. The total training epochs are set as 100 and the batch size is 8 by default. The learning rate is $2{\rm e}^{-4}$ initially, and dropped by $0.1$ after $[48,57,88]$ epochs. The evaluation is run with 3 random seeds. Code will be available online after paper acceptance.
@@
Table 1: Anomaly detection and localization results on MVTecAD. All methods are evaluated under the unified case. /normal-⋅normal-$\cdot$/$\cdot$ / ⋅ means the image-level and pixel-level AUROCs.
@@ 4.3 Ablation Studies
Hierarchical Gaussian Mixture Modeling. 1) Tab. 3a verifies our confirmation that mapping to multiple class centers is effective. With the fixed multiple centers (FMC), image-level and pixel-level AUROCs can be improved by 4.1% and 1.8%, respectively. By employing the inter-class Gaussian mixture modeling to learn the latent multi-class distribution, the AUROCs can further be improved by 1.4% and 0.9%. 2) The effectiveness of Mutual information maximization (MIM) is proven in Tab. 3a, where adding MIM brings promotion by 1.8% for detection. This shows that to better learn the complex multi-class distribution, it is necessary to endow the model class discrimination ability to avoid multiple centers collapsing into the same center. 3) Tab. 3a confirms the efficacy of intra-class mixed class centers learning. With the FMC as the baseline, introducing to learn intra-class mixed class centers could bring an increase of 1.6% for detection and 1.0% for localization, respectively. Finally, combining these, we form the hierarchical Gaussian mixture modeling method to achieve the best results.

### v2

@@ Abstract
Unified anomaly detection (AD) is one of the most valuable challenges for anomaly detection, where one unified model is trained with normal samples from multiple classes with the objective to detect anomalies in these classes. For such a challenging task, popular normalizing flow (NF) based AD methods may fall into a “homogeneous mapping” issue, where the NF-based AD models are biased to generate similar latent representations for both normal and abnormal features, and thereby lead to a high missing rate of anomalies. In this paper, we propose a novel Hierarchical Gaussian mixture normalizing flow modeling method for accomplishing unified Anomaly Detection, which we call HGAD. Our HGAD consists of two key components: inter-class Gaussian mixture modeling and intra-class mixed class centers learning. Compared to the previous NF-based AD methods, the hierarchical Gaussian mixture modeling approach can bring stronger representation capability to the latent space of normalizing flows. In this way, we can avoid mapping different class distributions into the same single Gaussian prior, thus effectively avoiding or mitigating the “homogeneous mapping” issue. We further indicate that the more distinguishable different class centers, the more conducive to avoiding the bias issue. Thus, we further propose a mutual information maximization loss for better structuring the latent feature space. We evaluate our method on four real-world AD benchmarks, where we can significantly improve the previous NF-based AD methods and also outperform the SOTA unified AD methods. The code will be available at https://github.com/xcyao00/HGAD.
@@ 1 Introduction
Anomaly detection has received increasingly wide attentions and applications in different scenarios, such as industrial defect detection [7, 23, 42, 43, 13, 39], video surveillance [1, 31], medical lesion detection [33, 46], and road anomaly detection [37, 10]. Considering the highly scarce anomalies and diverse normal classes, most previous AD studies have mainly devoted to unsupervised one-class learning, i.e., learning one specific AD model by only utilizing one-class normal samples and then detecting anomalies in this class111 Class means the category of the object in the image, for industrial AD, it refers to the industrial product category, e.g., in Fig. 1, cable, carpet, etc.. However, such a one-for-one paradigm would require more human labor, time, and computation costs when training and testing on many product categories, and also underperform when the one normal class has large intra-class diversity.
@@
In this work, we aim to tackle a more practical task: unified anomaly detection. As shown in Fig. 1b, one unified model is trained with normal samples from multiple classes, and the objective is to detect anomalies for all these classes without any fine-tuning. Nonetheless, solving such a task is quite challenging. Currently, there are two reconstruction-based AD methods for tackling the challenging unified AD task, UniAD [42] and PMAD [41]. But the reconstruction-based methods may fall into the “identical shortcut reconstruction” dilemma [42], where anomalies can also be well reconstructed, resulting in the failure of anomaly detection. UniAD and PMAD attempt to mask the adjacent or suspicious anomalies to avoid identical reconstruction. However, due to the diverse scale and shape of anomalies, the masking mechanism cannot completely avoid the abnormal information leakage during reconstruction, the risk of identical reconstruction is still existing. To this end, we consider designing the unified AD model from the normal data distribution learning perspective. The advantage is that we will no longer face the abnormal information leakage risk in principle, as our method is based on normal data distribution and radically avoids reconstruction. Specifically, we employ normalizing flows (NF) to learn the normal data distribution [15].
However, we find that the NF-based AD methods perform unsatisfactorily when applied to the unified AD task. They usually fall into a “homogeneous mapping” issue (see Sec. 3.2), where the NF-based AD models are biased to generate large log-likelihoods for both normal and abnormal inputs (see Fig. 2b). We explain this issue as: the NF-based AD methods [15, 43] employ the uni-modal Gaussian prior to learn the multi-class distribution (multi-modal). This can be seen as learning a mapping from a heterogeneous space to the latent homogeneous space. To learn the mapping well, the network may be prompted to take a bias to concentrate on the coarse-grained common characteristics (e.g., local pixel correlations) and suppress the fine-grained distinguishable characteristics (e.g., semantic content) among different class features [17]. Consequently, the network homogeneously maps different class features to the close latent embeddings. Thus, even anomalies can obtain large log-likelihoods and become less distinguishable.
@@
1. We propose a novel unified AD method: HGAD, where the hierarchical Gaussian mixture modeling approach can bring stronger representation capability to the latent space of normalizing flows for accomplishing the unified AD task.
@@ 2 Related Work
Anomaly Detection. 1) Reconstruction-based approaches are the most popular AD methods. These methods rely on the assumption that models trained by normal samples would fail in abnormal image regions. Many previous works attempt to train AutoEncoders [21, 44], Variational AutoEncoders [19] and GANs [29, 2] to reconstruct the input images. However, these methods face the “identical shortcut” problem [42]. 2) Embedding-based approaches recently show better AD performance by using ImageNet pre-trained networks as feature extractors [6, 11]. PaDiM [12] extract pre-trained features to model Multivariate Gaussian distribution for normal samples, then utilize Mahalanobis distance to measure the anomaly scores. PatchCore [23] extends on this line by utilizing locally aggregated features and introducing greedy coreset subsampling to form nominal feature banks. 3) Knowledge distillation assumes that the student trained to learn the teacher on normal samples could only regress normal features but fail in abnormal features [8]. Recent works mainly focus on feature pyramid [27, 38], reverse distillation [13], and asymmetric distillation [26]. 4) Unified AD approaches attempt to train a unified AD model to accomplish anomaly detection for multiple classes. UniAD [42], PMAD [41] and OmniAL [48] are three existing methods in this new direction. UniAD is a transformer-based reconstruction model with three improvements, it can perform well under the unified case by addressing the “identical shortcut” issue. PMAD is a MAE-based patch-level reconstruction model, which can learn a contextual inference relationship within one image rather than the class-dependent reconstruction mode. OmniAL is a unified CNN framework with anomaly synthesis, reconstruction and localization improvements.
@@
Normalizing Flows in Anomaly Detection. In anomaly detection, normalizing flows are employed to learn the normal data distribution [24, 15, 43, 40, 18], which maximize the log-likelihoods of normal samples during training. Rudolph et al. [24] first employ NFs for anomaly detection by estimating the distribution of pre-trained features. In CFLOW-AD [15], the authors further construct NFs on multi-scale feature maps to achieve anomaly localization. Recently, fully convolutional normalizing flows [25, 43] have been proposed to improve the accuracy and efficiency of anomaly detection. In BGAD [40], the authors propose a NF-based AD model to tackle the supervised AD task. In this paper, we mainly propose a novel NF-based AD model (HGAD) with three improvements to achieve much better unified AD performance.
@@ 3.2 Revisiting Normalizing Flow Based Anomaly Detection Methods
Figure 2: Comparison between many-to-one and our unified NF-based AD methods on MVTecAD. (a) and (c) show the training losses and the testing anomaly detection and localization AUROCs. (b) shows that the many-to-one NF-based AD model (i.e., training the one-for-one AD method, FastFlow, on multiple classes simultaneously) may have an obvious normal-abnormal overlap under the unified case, while ours (d) can bring better normal-abnormal distinguishability.
@@ 3.3 Hierarchical Gaussian Mixture Normalizing Flow Modeling
Figure 3: Model overview. The extracted feature vectors are sent into the normalizing flow model for transforming into latent embeddings. Following [15], we add positional embeddings (i.e., sinusoidal position encoding) to each invertible layer as they are effective for NF-based AD methods. Then, we employ our hierarchical Gaussian mixture modeling approach to fit the latent embeddings, which can assist the model against learning the “homogeneous mapping”.
@@ 3.4 Anomaly Scoring
As shown in Alg. 1, we actually extract K-level feature maps and construct the NF model at each level. For each test input feature $x^{k}$ from level-$k$, $k\in K$, we can calculate its intra-class log-likelihood ${\rm log}p_{\theta}(x^{k})=\mathop{{\rm logsumexp}}_{i}(-||\varphi_{\theta}(x^{k})-\mu_{i}^{y}||^{2}_{2}/2+c_{i}^{y})+{\rm log}|{\rm det}J|-d/2{\rm log}(2\pi)$ and inter-class negative entropy $nh(x^{k})=\sum_{y}\mathop{{\rm softmax}}_{y}(-||\varphi_{\theta}(x^{k})-\mu_{1}^{y^{\prime}}||^{2}_{2}/2)\cdot\mathop{{\rm logsoftmax}}_{y}(-||\varphi_{\theta}(x^{k})-\mu_{1}^{y^{\prime}}||^{2}_{2}/2)$. Note that $y$ in $c_{i}^{y}$ denotes which class $x^{k}$ belongs to, not whether $x^{k}$ is normal or abnormal. Next, we convert the log-likelihood to likelihood $p_{\theta}(x^{k})={\rm e}^{{\rm log}p_{\theta}(x^{k})}$. Then, we upsample all $p_{\theta}(x^{k})$ in the level-$k$ to the input image resolution $(H\times W)$ using bilinear interpolation $P_{k}=b(p_{\theta}(x^{k}))\in\mathbb{R}^{H\times W}$. Finally, we calculate anomaly score map $S_{l}$ by aggregating all upsampled likelihoods as $S_{l}=1-\sum_{k=1}^{K}P_{k}$ (see Sec. 3.1). For the inter-class negative entropy, we also follow the above steps to convert to anomaly score map $S_{e}$. Then the final anomaly score map is obtained by combining the two maps $S=S_{l}\odot S_{e}$, where the $\odot$ is the element-wise multiplication. In this way, even if anomalies fall into the inter-class Gaussian mixture distribution, they are usually in the low-density regions among the inter-class class centers. So, we can still ensure that anomalies are out-of-distribution through intra-class log-likelihoods (please see App. 0.A.4 for more discussions).
@@ 3.5 Further Discussions
In App. 0.A.3, we further provide more discussions with the “identical shortcut” issue and point out that the “homogeneous mapping” is not intrinsically equal to the “identical shortcut” issue. Compared to UniAD, our method is an effective exploration for unified anomaly detection in the direction of normalizing flow based anomaly detection. In App. 0.A, we also further discuss the limitations, applications, effective guarantee, and complexity of our method. We also provide an information-theoretic view in App. 0.F.
@@ 4.2 Main Results
Setup. For each dataset, we train one unified model with images from different classes. We use Efficient-b6 [32] as the feature extractor. The parameters of the feature extractor are frozen during training. The layer numbers of the NF models are all 12. The number of inter-class centers is always equal to the number of classes in the dataset. The number of intra-class centers is set as $10$ for all datasets (see ablation study in Sec. 4.3). We use the Adam [22] optimizer with weight decay $1{\rm e}^{-4}$ to train the model. The total training epochs are set as 100 and the batch size is 8 by default. The learning rate is $2{\rm e}^{-4}$ initially, and dropped by $0.1$ after $[48,57,88]$ epochs. The evaluation is run with 3 random seeds.
@@
Table 1: Anomaly detection and localization results on MVTecAD. All methods are evaluated under the unified case. $\cdot$/$\cdot$ / ⋅ means the image-level and pixel-level AUROCs.
@@ 4.3 Ablation Studies
Hierarchical Gaussian Mixture Modeling. 1) Tab. 3a verifies our confirmation that mapping to multiple class centers is effective. With the fixed multiple centers (FMC, which is only based on fixed multiple centers but without Gaussian mixture modeling), image-level and pixel-level AUROCs can be improved by 4.1% and 1.8%, respectively. By employing the inter-class Gaussian mixture modeling to learn the latent multi-class distribution, the AUROCs can further be improved by 1.4% and 0.9%. 2) The effectiveness of Mutual information maximization (MIM) is proven in Tab. 3a, where adding MIM brings promotion by 1.8% for detection. This shows that to better learn the complex multi-class distribution, it is necessary to endow the model class discrimination ability to avoid multiple centers collapsing into the same center. 3) Tab. 3a confirms the efficacy of intra-class mixed class centers learning. With the FMC as the baseline, introducing to learn intra-class mixed class centers could bring an increase of 1.6% for detection and 1.0% for localization, respectively. Finally, combining these, we form the hierarchical Gaussian mixture modeling method to achieve the best results.
@@
Optimization Strategy. We found that simultaneously optimizing inter-class and intra-class centers by the beginning would bring instability. Thus, we employ a two-stage optimization strategy (see App. 0.C for details). The strategy can better decouple the inter-class and intra-class learning processes, thereby bringing better results.